# Classical Benchmarks for Variational Quantum Eigensolver Simulations of the Hubbard Model

### v1

@@ III.5 Scaling up and extracting correlation functions
Overall, our work so far has focused on the quantitative benchmarking of VQE simulations of the Hubbard model. However, qualitative features of the model can be important to reproduce, and indeed VQAs have been shown in various examples to yield ground states that display the well-known antiferromagnetic behavior of the Hubbard model in the strong coupling regime [40, 31]. Our VQE implementation based on tensor networks can recover such behaviors; in Fig. 8, we visualize the spin correlation function
$\langle C_{ij}\rangle=\langle S_{i}^{z}S_{j}^{z}\rangle-\langle S_{i}^{z}\rangle\langle S_{j}^{z}\rangle$ (16)
in the case of $U/t=8$ for a $1\times 6$ and $1\times 12$ chain, within VQE based on energy- and overlap-based optimizations, and within DMRG. To obtain the VQE results here we have used 13 layers of the NP ansatz. We observe that for the smaller chain, the spin properties from both VQE approaches and DMRG are in near perfect agreement. For the larger $1\times 12$ chain, VQE based on an energy minimization fails to fully capture longer-range spin correlations, in agreement with previous findings [39, 40]. However, the VQE ground state obtained through overlap-based optimization is in near-perfect agreement to the DMRG one in terms of its long-range spin correlation, highlighting the utility of this approach for computing such features.
IV Conclusions and outlook
@@
In this work we have presented a rigorous benchmark of the quantitative features of the solution of the Hubbard model using variational quantum eigensolvers. We find that the NP ansatz of Ref. [35] yields energies with the smallest error relative to a DMRG reference, compared to the EP and UCC ansätze. However, even when using the NP ansatz, the error in the ground state energy increases for larger lattices, and ultimately plateaus with increasing number of optimization parameters. Moreover, we have found that systems with stronger electronic correlations pose an even greater challenge in terms of their accurate quantitative description, and require more optimization parameters. At the same time, spatial inhomogeneity in the parameters of the Hubbard model does not seem to strongly affect the convergence rate of the results, nor does the addition of nearest-neighbor Coulomb interactions. This is encouraging as we are moving towards the description of more complex versions of the Hubbard model, representing real materials through schemes such as ab initio downfolding. We have also shown, in agreement with previous work, that the qualitative characteristics of the ground state wavefunction of the one-dimensional Hubbard model reproduce the expected magnetic behavior depending on the $U/t$ ratio. We found that while for smaller lattices the VQE solutions obtained via energy minimization can give quantitatively accurate results for spin correlation functions, for larger systems it struggles to capture long-range correlations. However, a VQE optimization based on maximizing the overlap to the solution of a classical reference method succeeds in describing these long-range phenomena. While of course ultimately it is the goal of the community to utilize quantum computers to obtain the properties of Hubbard model ground states for systems beyond the reach of classical methods, where computing the overlap to the “true” ground state will not be possible, this result still highlights the potential of an overlap-based optimization using an approximate classical reference state to provide a strong starting point for energy-based VQE. The overlap-based optimization results also highlight that the NP ansatz is expressive enough to capture features like long-range correlations, suggesting that improvements to the energy-based optimization approach may be able to remedy these issues.
@@
Moving forward, it will be important to extend the applicability of current approaches to the solution of Hubbard models with multiple electronic bands, as these multi-band models describe several physical phenomena of relevance to applications, including superconductivity and excitonic insulating behavior [19, 58, 59, 60]. Given the challenge with systematically converging to the global energy minimum with current variational approaches as demonstrated here, even in the one-band case and particularly for the strongly correlated regime, it will be important to develop flexible approaches to achieve this. One possibility is to explore other ansatz generation schemes, such as adaptive ansätze [61, 62, 63, 64, 43, 65] or problem-inspired ansätze tailored for the strongly-correlated regime.
V Acknowledgments
@@
The authors acknowledge useful discussions with Bryan Clark and Erik Gustafson. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359. We are grateful for support from NASA Ames Research Center. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award ASCR-ERCAP0024469.

### v2

@@ 3.4 Extracting correlation functions and overlap-based optimization
Our work so far has focused on the quantitative benchmarking of VQE simulations of the Hubbard model in terms of predicting the ground state energy. However, qualitative features of the model can be important to reproduce, and indeed VQAs have been shown in various examples to yield ground states that display the well-known antiferromagnetic behavior of the Hubbard model, which is most pronounced in the strong coupling regime [46, 33]. To understand the performance of VQE in terms of recovering such behaviors, we compute the spin correlation function
$\langle C_{ij}\rangle=\langle S_{i}^{z}S_{j}^{z}\rangle-\langle S_{i}^{z}\rangle\langle S_{j}^{z}\rangle,$ (17)
[…]
Figure 9: Comparison of the fidelity of the VQE ground state with respect to DMRG, as obtained from an energy-based optimization (solid lines) and an overlap-based optimization (dashed lines), when using the NP ansatz, and for 2D lattices with $U/t=2$ (panel a) and $U/t=8$ (panel b). Comparison of the energy of the VQE ground state with respect to DMRG, as obtained from an energy-based optimization (solid lines) and an overlap-based optimization (dashed lines), when using the NP ansatz, and for 2D lattices with $U/t=2$ (panel c) and $U/t=8$ (panel d). The shaded regions indicate the full range of values found by optimizing ten independent configurations in each case.
The improved description of long-range correlation functions from an overlap-based VQE approach is because of the higher fidelity of the ground state wavefunctions produced by this approach. In order to systematically benchmark this behavior, we visualize in Figs. 9a and 9b the fidelity $\mathcal{F}=|\bra{\Psi_{\rm VQE}}\ket{\Psi_{\rm DMRG}}|^{2}$ of the VQE ground state from the energy- (solid lines) and overlap-based (dashed-lines) approaches, when using the NP ansatz for 2D lattices, and for $U/t=2$ and $U/t=8$. We see that with the addition of more parameters to the optimization, both methods converge towards the limit of $\mathcal{F}=1$, except for the largest system size $3\times 3$. However, the convergence of the overlap-based optimization is generally faster, especially for larger lattices, a feature which is particularly prominent in the case of $U/t=8$, where strong electronic correlations are present. The fact that the overlap-based optimization performs better in terms of yielding higher-fidelity ground states is not surprising, given that this method is tailored specifically to minimize the loss function of Eq. (5), i.e., to minimize the infidelity of the VQE ground state with respect to the DMRG one. It is worth emphasizing that the higher fidelity solutions produced via overlap-based optimization do not always possess lower energies compared to the lower-fidelity wavefunctions obtained via energy-based minimization, and their energy error can in fact often be even somewhat greater than that from energy-based optimization. The VQE energies obtained with the two approaches are visualized in Figs. 9c and 9d for $U/t=2$ and $U/t=8$ respectively, for $2\times 3$ and $3\times 3$ lattices, with all the ground-state energy values of the overlap-based approach given in the Supplemental Material. The higher energies that are produced by the overlap-based optimization are not surprising given the nature of this method. If we consider the solution to be of the form |$\ket{\Psi}_{VQE}=\alpha\ket{\Psi}_{GS}+\beta\ket{\chi}$, where |$\ket{\Psi}_{GS}$ the true ground state and |$\ket{\chi}$ some arbitrary wavefunction, the aim of the overlap-based approach is to produce solutions which maximize the value of $|\alpha|$. However, depending on the nature of the admixture $\beta\ket{\chi}$, for the energy of which there are no guarantees, the state |$\ket{\Psi}$ could assume a range of energies higher than that achieved through an energy-based minimization. Nevertheless, we see that while the minimum energy produced by the overlap-based approach can sometimes be higher than that of the energy-based approach, in practice the two results are often within the margin of error of each other. Therefore, given the superior ground state fidelity produced by an overlap-based optimization, this approach can provide a balanced approach to finding the ground state of the Hubbard model. It also suggests using overlap-optimized wavefunctions as input for further energy minimization, which is feasible even for larger lattices when the DMRG wavefunctions are no longer exact.
4 Conclusions and outlook
@@
In this work we have presented a rigorous benchmark of the quantitative features of the solution of the Hubbard model using variational quantum eigensolvers. We find that the NP ansatz of Ref. [41] yields energies with the smallest error relative to a DMRG reference, compared to the EP and UCC ansätze. However, even when using the NP ansatz, the error in the ground state energy increases for larger lattices, and ultimately plateaus with increasing number of parameters, due to the optimization becoming stuck in local minima. While our exact calculations on classical computers are not limited by barren plateau phenomena, these will make the optimization even more challenging on quantum hardware. Moreover, we have found that systems with stronger electronic correlations pose an even greater challenge in terms of their accurate quantitative description, and require more optimization parameters. At the same time, spatial inhomogeneity in the parameters (i.e., disorder) does not seem to strongly affect the convergence rate of the results, nor does the addition of nearest-neighbor Coulomb interactions. This is encouraging as we are moving towards the description of more complex versions of the Hubbard model, representing real materials through schemes such as ab initio downfolding [25, 50]. We expect our results to serve as an important reference for future applications of VQE for the solution of Hubbard models and strongly-correlated materials, providing an understanding of the accuracy and limitations that can be expected depending on the features of a given system.
We have also shown, in agreement with previous work, that the qualitative characteristics of the ground state wavefunction of the one-dimensional Hubbard model reproduce the expected magnetic behavior depending on the $U/t$ ratio. We found that while for smaller lattices the VQE solutions obtained via energy minimization can give quantitatively accurate results for spin correlation functions, for larger systems it struggles to capture long-range correlations. However, a VQE optimization based on maximizing the overlap to the solution of a classical reference method succeeds in describing these long-range phenomena. While of course ultimately it is the goal of the community to utilize quantum computers to obtain the properties of Hubbard model ground states for systems beyond the reach of classical methods, where computing the overlap to the “true” ground state will not be possible, this result still highlights the potential of an overlap-based optimization using an approximate classical reference state to provide a strong starting point for energy-based VQE. The overlap-based optimization results also highlight that the NP ansatz is expressive enough to capture features like long-range correlations.
@@
Moving forward, it will be important to extend the applicability of current approaches to the solution of Hubbard models with multiple electronic bands, as these multi-band models describe several physical phenomena of relevance to applications, including superconductivity and excitonic insulating behavior [19, 72, 73, 74]. Given the challenge with systematically converging to the global energy minimum with current variational approaches as demonstrated here, even in the one-band case and particularly for the strongly correlated regime, it will be important to develop flexible approaches to achieve this. One possibility is to explore other ansatz generation schemes, such as adaptive ansätze [75, 76, 77, 78, 49, 79] or problem-inspired ansätze tailored for the strongly-correlated regime. Alternatively, optimized variational wavefunctions that have converged to local minima can be used as initial states of a subsequent quantum subspace expansion, which is known to converge faster for a larger overlap between the initial state and the true ground state [80, 81, 82, 83].
5 Acknowledgments
@@
The authors acknowledge useful discussions with Bryan Clark and Erik Gustafson. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359. We are grateful for support from NASA Ames Research Center. Ames National Laboratory is operated for the U.S. Department of Energy by Iowa State University under Contract No. DE-AC02-07CH11358. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC awards HEP-ERCAP0029167 and DDR-ERCAP0029710.